# Learning to Play Two-Player Perfect-Information Games without Knowledge

### v4

@@ 2.3 Action Selection Distribution
One of the problems related to reinforcement learning is the exploration-exploitation dilemma [9]. It consists of choosing between exploring new states to learn new knowledge and exploiting the acquired knowledge. Many techniques have been proposed to deal with this dilemma [41]. However, most of these techniques do not scale because their application requires memorizing all the encountered states. For this reason, in the context of games with large numbers of states, some approaches use probabilistic exploration [36, 2, 9, 42]. With this approach, to exploit is to play the best action and to explore is to play uniformly at random. More precisely, a parametric probability distribution is used to associate with each action its probability of being played. The parameter associated with the distribution corresponds to the exploration rate (between 00 and $1$), which we denote $\epsilon$ (the exploitation rate is therefore $1-\epsilon$, which we denote $\epsilon^{\prime}$). The rate is often experimentally fixed. Simulated annealing [43] can, however, be applied to avoid choosing a value for this parameter. In this case, at the beginning of reinforcement learning, the parameter is $1$ (we are just exploring). It gradually decreases until reaching 00 at the end of learning. The simplest action selection distribution is $\epsilon$-greedy [36] (of parameter $\epsilon$). With this distribution, the action is chosen uniformly with probability $\epsilon$ and the best action is chosen with probability $1-\epsilon$ (see also Algorithm 2).
Function $\epsilon$_greedy($s$, $v^{\prime}$)
      if probability $\frac{t}{t_{\max}}$  then
@@
                  return $\operatorname*{\mathrm{arg\,max}}_{a\in\text{{{actions(}}{\emph{\text{s}}}{{)}}}}v^{\prime}\left(s,a\right)$
@@
                  return $\operatorname*{\mathrm{arg\,min}}_{a\in\text{{{actions(}}{\emph{\text{s}}}{{)}}}}v^{\prime}\left(s,a\right)$
@@
Algorithm 2 $\epsilon$-greedy algorithm with simulated annealing used in the experiments of this article (see Table 1 for the definitions of symbols).
The $\epsilon$-greedy distribution has the disadvantage of not differentiating the actions (except the best action) in terms of probabilities. Another distribution is often used, correcting this disadvantage. This is the softmax distribution [42, 9]. It is defined by $P\left(a_{i}\right)=\frac{e^{\nicefrac{{v^{\prime}(s,a_{i})}}{{\tau}}}}{\sum_{j=1}^{n}e^{\nicefrac{{v^{\prime}(s,a_{j})}}{{\tau}}}}$ with $n$ the number of children of the current state $s$, $P\left(a_{i}\right)$ the probability of playing the action $a_{i}$ , $v^{\prime}(s,a_{i})$ the value of the state obtained after playing $a_{i}$ in $s$, $i\in\{0,\ldots,n-1\}$, and $\tau\in]0,+\infty[$ a parameter called temperature ($\tau\simeq 0$ : exploitation, $\tau\simeq+\infty$ : exploration).
2.4 Game of Hex
The game of Hex [44] is a two-player combinatorial strategy game. It is played on an empty $n\times n$ hexagonal board. We say that a $n\times n$ board is of size $n$. The board can be of any size, although the classic sizes are $11$, $13$ and $19$. In turn, each player places a stone of his color on an empty cell (each stone is identical). The goal of the game is to be the first to connect the two opposite sides of the board corresponding to its color. Figure 1 illustrates an end game. Although these rules are simplistic, Hex tactics and strategies are complex. The number of states and the number of actions per state are very large, similar to the game of Go. From the board size $11$, the number of states is, for example, higher than that of chess (Table $6$ of [45]). For any board size, the first player has a winning strategy [46] which is unknown, except for board sizes smaller than or equal to $10$ [47] (the game is weaky solved up to the size $10$). In fact, resolving a particular state is PSPACE-complete [48, 49]. There is a variant of Hex using a swap rule. With this variant, the second player can play in first action a special action, called swap, which swaps the color of the two players (i.e. they swap their pieces and their sides). This rule reduces the imbalance between the two players (without the swap rule, the first player has a very strong advantage). It is generally used in competitions.
Figure 1: A Hex end game of size 11 (white wins)
2.5 Hex Programs
Programs which learn the evaluation function by reinforcement have also been designed. These programs are NeuroHex [36], EZO-CNN [56], DeepEzo [57] and ExIt [20]. They learn from self-play. Unlike the other three programs, NeuroHex performs supervised learning (of a common Hex heuristic) followed by reinforcement learning. NeuroHex also starts its games with a state from a database of games. EZO-CNN and DeepEzo use knowledge to learn winning strategies in some states. DeepEzo also uses knowledge during confrontations. ExIt learns a policy in addition to the value of states and it is based on MCTS. It is the only program to have learned to play Hex without using knowledge. This result is, however, limited to the board size $9$. A comparison of the main characteristics of these different programs is presented in Table 2.
Programs Size Search Learning Network Use
Mohex-CNN 13 MCTS supervised convolutional policy
Mohex-3HNN 13 MCTS supervised residual policy, state, action
NeuroHex 13 none supervised, reinforcement convolutional state
EZO-CNN 7, 9, 11 Minimax reinforcement convolutional state
DeepEZO 13 Minimax reinforcement convolutional policy, state
ExIt 9 MCTS reinforcement convolutional policy, state
Table 2: Comparison of the main features of the latest Hex programs. These characteristics are respectively the board sizes on which learning is based, the used tree search algorithm, the type of learning, the type of neural network and its use (to approximate the values of states, actions and/or policy.
@@ 2.6 Games of Paper Experiments
We now briefly present the other games on which experiments are performed in this article, namely: Surakarta, Othello, Outer Open Gomoku, Clobber, Breakthrough, Amazons, Lines of Action, Santorini. They are all board games. All of these games (except Santorini so far) are present and recurring at the Computer Olympiads, the worldwide multi-games event in which computer programs compete against each other. Moreover, all these games (and their rules) are included (and available for free) in Ludii [58], a general game system.
@@ 2.6.1 Surakarta
Surakarta is a move and capture game (like the checkers game), the object of the game being to take all the opposing pieces. In his turn, a player can either move a piece to an empty square at a distance of $1$ or move a piece to a square occupied by an opponent’s piece under certain conditions and according to a mechanism specific to Surakarta (based on a movement circuit dedicated only to capture), allowing "long distance" capture.
2.6.2 Othello
Othello (also called Reversi) is a linear territory and encirclement game whose goal is to have more pieces than your opponent. In his turn, a player places a piece of his color on the board (only if he can make an encirclement, otherwise he pass his turn). There is an encirclement if an opponent’s line of pieces has at its two ends the piece that has just been placed and another piece of the player performing the encirclement. As a result of this encirclement, the encircled pieces are replaced by pieces from that player.
@@ 2.6.3 Outer Open Gomoku
Outer Open Gomoku is an alignment game. The object of the game is to line up at least $5$ pieces of its color. On his turn, a player places a piece of his color. In the first turn, the first player can only place a piece at a distance of $2$ from the sides of the board.
@@ 2.6.5 Breakthrough
Breakthrough is a move and capture game. The object of the game is to be the first to make one of his pieces reach the other side of the board. A piece can only move by moving forward one square (straight or diagonal). A capture can only be done diagonally.
@@ 2.6.6 Amazons
Amazons is a move and blocking game. In turn, a player moves one of his pieces in a straight line in any direction (like the queen of the game of Chess). Then he places a neutral piece (blocking movements like the players’ pieces) in any direction starting from the new position of the piece just moved (in the manner of the queen of the game of Chess). The goal of the game is to be the last to play.
@@ 2.6.7 Lines of Action
Lines of Action is a game of movement and regrouping. On his turn, a player can move one of his pieces in one direction as many squares as there are pieces in that direction. A piece cannot move if there is an opponent’s piece in its path, unless it is the square to arrive (in which case a capture is made). The goal is to have all of its pieces connected (at the same time).
@@ 2.6.8 Santorini
Santorini is a three-dimensional building and moving game. The goal of the game is to reach the 3rd floor of a building. In his turn, a player moves one of his pieces by one square then places the first floor on an adjacent empty square or increases a pre-existing construction by one floor (on which no player’s piece is located). A piece cannot move to a square having strictly more than one floor more than the square where it is located (a piece only go up one floor at a time and can descend as many floors as wanted). A move cannot be made to a square with 4 floors. A construction cannot be done on a square of 4 floors. A player who cannot play loses. Advanced mode (i.e. the use of power cards) is not used in the experiments in this article.
@@ 3 Data Use in Game Learning
In this section, we adapt and study tree learning (see Section 2.2) in the context of reinforcement learning and the use of non-linear adaptive evaluation functions. For this, we compare it to root learning and terminal learning in this context. We start by adapting tree learning, root learning, and terminal learning. Next, we describe the experiment protocol common to several sections of this article. Finally, we expose the comparison of tree learning with root learning and terminal learning.
@@ 3.1 Tree Learning
Function tree_learning($t_{\max}$)
@@
                  $S,\,T\leftarrow$search($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
@@
Algorithm 3 Tree learning (tree bootstrapping) algorithm (see Table 1 for the definitions of symbols). In this context, $S$ is the set of states which are non-leaves or terminal.
Function root_learning($t_{\max}$)
@@
                  $S,\,T\leftarrow$search($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
@@
Function terminal_learning($t_{\max}$)
@@
                  $S,\,T\leftarrow$search($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
@@
      Split $D$ in $m$ disjoint sets, denoted $\left\{D_{i}\right\}_{i=1}^{m}$, such that $D=\bigcup_{i=1}^{m}D_{i}$ and $\left|D_{i}\right|=B$ for each $i\in\left\{1,\ldots,m-1\right\}$
@@ 3.2 Common Experiment Protocol
The experiments of several sections share the same protocol. It is presented in this section. The protocol is used to compare different variants of reinforcement learning algorithms. A variant corresponds to a certain combination of elementary algorithms. More specifically, a combination consists of the association of a search algorithm (iterative deepening alpha-beta (with move ordering), MCTS (UCT with $c=0.4$ as exploration constant), $\mathrm{UBFM}$, …), of an action selection method ($\epsilon$-greedy distribution (used by default), softmax distribution, …), a terminal evaluation function $f_{\mathrm{t}}$ (the classic game gain (used by default), …), and a procedure for selecting the data to be learned (root learning, tree learning, or terminal learning). The protocol consists in carrying out a reinforcement learning of $48$ hours for each variant. At several stages of the learning process, matches are performed using the adaptive evaluation functions obtained by the different variants. Each variant is then characterized by a winning percentage at each stage of the reinforcement learning process. More formally, we denote by $f_{\theta_{h}}^{c}$ the evaluation generated by the combination $c$ at the hour $h$. Each combination is evaluated every hour by a winning percentage. The winning percentage of a combination $c$ at a hour $h\leq 48$ (i.e. of $f_{\theta_{h}}^{c}$) is computed from matches against each combination $c^{\prime}$ at final time $h=48$, i.e. against each $f_{\theta_{48}}^{c^{\prime}}$ (there is one match in first player and another in second player per pair of combination). The matches are made by using alpha-beta at depth $1$.
@@
In addition to the curve, the different variants are also compared in relation to their final winning percentage, i.e. at the end of the learning process. Unlike the experiment of the evolution of winning percentages, in the comparison of the different variants at the final stage, each evaluation $f_{\theta_{48}}^{c}$ confronts each other evaluation $f_{\theta_{48}}^{c^{\prime}}$ of all the repetitions. In other words, this experiment consists of performing an all-play-all tournament with all the evaluation functions generated during the different repetitions. The presented winning percentage of a combination is still the average over the repetitions. The matches are also made by using alpha-beta at depth $1$. These percentages are shown in tables.
@@
All the experiments involving MCTS were also performed with $c=\sqrt{2}$ as exploration constant. The results are similar.

### v5

@@ 2.3 Action Selection Distribution
One of the problems related to reinforcement learning is the exploration-exploitation dilemma [7]. It consists of choosing between exploring new states to learn new knowledge and exploiting the acquired knowledge. Many techniques have been proposed to deal with this dilemma [43]. However, most of these techniques do not scale because their application requires memorizing all the encountered states. For this reason, in the context of games with large numbers of states, some approaches use probabilistic exploration [38, 2, 7, 44]. With this approach, to exploit is to play the best action and to explore is to play uniformly at random. More precisely, a parametric probability distribution is used to associate with each action its probability of being played. The parameter associated with the distribution corresponds to the exploration rate (between 00 and $1$), which we denote $\epsilon$ (the exploitation rate is therefore $1-\epsilon$, which we denote $\epsilon^{\prime}$). The rate is often experimentally fixed. Simulated annealing [45] can, however, be applied to avoid choosing a value for the parameter. In that case, at the beginning of reinforcement learning, the parameter is $1$ (we are just exploring). It gradually decreases until reaching 00 at the end of the learning process. The simplest action selection distribution is $\epsilon$-greedy [38] (of parameter $\epsilon$). With the distribution $\epsilon$-greedy, the action is chosen uniformly with probability $\epsilon$ and the best action is chosen with probability $1-\epsilon$ (see also Algorithm 2).
Function $\epsilon$_greedy($s$, $v$)
      if  $\mathrm{random}\left(\right)\leq\frac{t}{t_{\max}}$  then
@@
                  return $\operatorname*{\mathrm{arg\,max}}_{a\in\text{{{actions(}}{\emph{\text{s}}}{{)}}}}v\left(s,a\right)$
@@
                  return $\operatorname*{\mathrm{arg\,min}}_{a\in\text{{{actions(}}{\emph{\text{s}}}{{)}}}}v\left(s,a\right)$
@@
Algorithm 2 $\epsilon$-greedy algorithm with simulated annealing ($\epsilon=1-\frac{t}{t_{\max}}$) used in the experiments of this article (see Table 1 for the definitions of symbols).
The $\epsilon$-greedy distribution has the disadvantage of not differentiating the actions (except the best action) in terms of probabilities. Another distribution is often used, correcting this disadvantage. This is the softmax distribution [44, 7]. It is defined by $P\left(a_{i}\right)=\frac{e^{\nicefrac{{v(s,a_{i})}}{{\tau}}}}{\sum_{j=1}^{n}e^{\nicefrac{{v(s,a_{j})}}{{\tau}}}}$ with $n$ the number of children of the current state $s$, $P\left(a_{i}\right)$ the probability of playing the action $a_{i}$ , $v(s,a_{i})$ the value of the state obtained after playing $a_{i}$ in the current state $s$, $i\in\{0,\ldots,n-1\}$, and $\tau\in]0,+\infty[$ a parameter called temperature ($\tau\simeq 0$ : exploitation, $\tau\simeq+\infty$ : exploration).
@@ 2.4 Games of Paper Experiments
We now briefly present the games on which experiments are performed in this article, namely: Hex, Arimaa, Morpion Solitaire, Othello, Surakarta, Outer Open Gomoku, Clobber, Breakthrough, Amazons, Lines of Action, and Santorini. They are all board games. All of these games are present and recurring (except Morpion Solitaire) at the Computer Olympiad, the worldwide multi-games event in which computer programs compete against each other. Moreover, all these games (and their rules) are included (and available for free) in Ludii [46], a general game system.
2.4.1 Hex
Hex is a two-player combinatorial strategy game played on an $n\times n$ hexagonal board. Each player takes turns placing a stone of their color on an empty cell. The objective is to be the first to connect the two opposite sides of the board corresponding to their color. A common variant of Hex includes the swap rule, where the second player can choose, on their first turn, to swap roles and sides with the first player. This rule helps balance the game by mitigating the first player’s advantage and is typically used in competitions. It is still used at the Computer Olympiad, and we use it in this paper.
2.4.2 Arimaa
Arimaa is a game with similarities with Chess. Each player has a different set of pieces (identical between the two players): one elephant, one camel, two horses, two dogs, two cats, and eight rabbits. The pieces types are fully ordered: elephant $>$ camel $>$ horse $>$ dog $>$ cat $>$ rabbit. A higher piece can move (push or pull) a lower piece. A piece which is orthogonally adjacent to a stronger opposing piece is frozen, unless it is also adjacent to a friendly piece. If a piece is on a trap square without being adjacent to a friendly piece, it is removed from the board. A player wins when one of its rabbits reaches the other side of the board. A player without rabbits loses. It is played on a $8\times 8$ grid like Chess (the cases $(2,2)$, $(2,5)$, $(5,2)$, $(5,5)$ are traps). At the start, the board is empty. The first player places its pieces on its side of the board then the second player does the same. On his turn, the player can make 4 moves of one square (with one or more pieces). Pushing or pulling an opposing piece counts as two moves. Details about rules are available in https://arimaa.com/arimaa/.
2.4.3 Morpion Solitaire
The game is played on a grid of unlimited size. The starting configuration includes a set of points already placed on the grid in the shape of a Greek cross whose side has 4 points (the cross is made up of a total of 36 points).
A move in this game consists of two steps. First, the player places a new point on the grid so as to form an alignment of five adjacent points horizontally, vertically or diagonally (which have not already been connected by a line). Then, he connects this alignment by drawing a line. An alignment cannot be placed as an extension of a previous alignment.
The aim of the game is to place as many points as possible before reaching a situation where no new points can be placed.
2.4.4 Othello
Othello (also called Reversi) is a territory and linear encirclement game whose goal is to have more pieces than its opponent. In his turn, a player places a piece of his color on the board (only if he can make an encirclement, otherwise he pass his turn). There is an encirclement if an opponent’s line of pieces has at one of its ends the piece that the current player has just placed, and has at the other end, another piece of the current player. As a result of this encirclement, the encircled opponent’s pieces are replaced by pieces from the current player.
@@ 2.4.5 Surakarta
Surakarta is a move and capture game (like draughts). The goal of the game is to take all the opposing pieces. In his turn, a player can either move a piece to an empty square at a distance of $1$ or move a piece to a square occupied by an opponent’s piece under certain conditions and according to a mechanism specific to Surakarta (based on a movement circuit dedicated only to capture), allowing “long distance” capture.
@@ 2.4.6 Outer Open Gomoku
Outer Open Gomoku is an alignment game. The goal of the game is to line up at least $5$ pieces of its color. On his turn, a player places a piece of his color. In the first turn, the first player can only place a piece at a distance of $2$ from the sides of the board.
@@ 2.4.8 Breakthrough
Breakthrough is a move and capture game. The goal of the game is to be the first to make one of his pieces reach the other side of the board. A piece can only move by moving forward one square (straight or diagonal). A capture can only be done diagonally.
@@ 2.4.9 Amazons
Amazons is a move and blocking game. In turn, a player moves one of his pieces in a straight line in any direction (like the queen of the game of Chess). Then he places a neutral piece in any direction starting from the new position of the piece just moved (always in the manner of the queen of the game of Chess ). Any piece (neutral or player-owned) blocks placements and movements. The goal of the game is to be the last to play.
@@ 2.4.10 Lines of Action
Lines of Action is a game of movement and regrouping. On his turn, a player can move one of his pieces in one direction as many squares as there are pieces in that direction. A piece cannot move if there is an opponent’s piece in its path, unless it is the square to arrive (in which case a capture is made). The goal is to have all of its pieces connected.
@@ 2.4.11 Santorini
Santorini is a three-dimensional building and moving game. The goal of the game is to reach the 3rd floor of a building. In his turn, a player moves one of his pieces by one square then places the first floor on an adjacent empty square or increases a pre-existing construction by one floor (on which no player’s piece is located). A piece cannot move to a square having strictly more than one floor more than the square where it is located (a piece only go up one floor at a time and can descend as many floors as wanted). A move cannot be made to a square with 4 floors. A construction cannot be done on a square of 4 floors. A player who cannot play loses. The advanced mode (i.e. the use of power cards) is not used in the experiments in this article.
@@ 3 Data Use in Game Learning
In this section, we adapt and study tree learning (see Section 2.2) in the context of reinforcement learning and the use of non-linear adaptive evaluation functions. For this purpose, we compare it to root learning and terminal learning. We start by adapting tree learning, root learning, and terminal learning. We also define in Section 8 a variant of experience replay that we call stratified experience replay. Next, we describe the experiment protocol common to several sections of this article. Finally, we expose the comparison of tree learning with root learning and terminal learning.
@@ 3.1 Tree Learning
Function tree_learning($t_{\max}$, $\tau$)
@@
                  $S,\,T\leftarrow$search($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$, $\tau$)
@@
Algorithm 3 Tree learning (tree bootstrapping) algorithm (see Table 1 for the definitions of symbols). For tree learning, $S$ is the set of states which are non-leaves or terminal.
Function root_learning($t_{\max}$, $\tau$)
@@
                  $S,\,T\leftarrow$search($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$, $\tau$)
@@
Function terminal_learning($t_{\max}$, $\tau$)
@@
                  $S,\,T\leftarrow$search($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$, $\tau$)
@@
      Split $D$ in $m$ disjoint sets, denoted $\left\{D_{i}\right\}_{i=1}^{m}$, such that $D=\bigcup_{i=1}^{m}D_{i}$ and $\left|D_{i}\right|=B$ for each $i\in\left\{1,\ldots,m\right\}$
@@
3.2 Stratified Experience Replay
In this section, we introduce the technique we call Stratified Experience Replay, which is a variation of Experience Replay. The targeted objective of Stratified Experience Replay is to reduce the variance of learning processes.
Stratified Experience Replay consists in performing a stratified sampling without replacement on the replay buffer of Experience Replay whose data has been grouped in relation to the matches that generated it. Stratified sampling imposes the same proportion of data for each previous match, so we have a good level of variability during each learning phase. The absence of replacement makes that no data is unnecessarily repeated and no data is forgotten.
Stratified Experience Replay depends on two parameters: $\mu$ the number of previous matches kept in memory and $\delta$ the duplication factor: the number of times a piece of data must be learned during the learning phases. The data is stored in two datasets, the data to be learned stored in $\mu$ sets $D_{i}$ with $i\in\{1,\ldots,\mu\}$, each corresponding to the data of the $i$-th previous match that must be learned and $\mu$ other sets $D^{\prime}_{i}$ with $i\in\{1,\ldots,\mu\}$, the archived data, which correspond to the data already learned from the $i$-th previous match. The set of data to learn for the new learning phase is then samples of each $D_{i}$ of size $\left\lceil\frac{\delta}{\mu}\left(\left|D_{i}\right|+\left|D^{\prime}_{i}\right|\right)\right\rceil$ (the total number of learned data is the number of generated data times the duplication factor $\delta$). The selected samples are removed from the $D_{i}$ and added to the $D^{\prime}_{i}$ The selected data is then separated into minibatches, so as to preserve the stratification as much as possible (the minibatches are thus in a way also stratified) and learning is then performed on each minibatch. The formalization of Stratified Experience Replay is described in Algorithm 8.
Function SGD_with_stratified_experience_replay($f_{\theta}$, $D$, $\mu$, $\delta$, $B$)
      append $\left(D,\emptyset\right)$ in $M$
      if $\left|M\right|>\mu$ then
            remove the oldest items of $M$ to have $\left|M\right|=\mu$
      $n\leftarrow\sum\left\{\left\lceil\frac{\delta}{\mu}\left(\left|D\right|+\left|D^{\prime}\right|\right)\right\rceil\,|\,\left(D,D^{\prime}\right)\in M\right\}$
      $g\leftarrow\max\left(1,\left\lfloor n/B\right\rfloor\right)$
      if $\left|\left\lfloor n/g\right\rfloor-B\right|>\left|\left\lfloor n/\left(g+1\right)\right\rfloor-B\right|$ then
            $g\leftarrow g+1$
      $G\leftarrow\left\{1,\ldots,g\right\}$
      foreach $i\in G$ do
            $S_{i}\leftarrow\emptyset$
      foreach $\left(D,D^{\prime}\right)\in M$ do
            $l\leftarrow\left\lceil\frac{\delta}{\mu}\left(\left|D\right|+\left|D^{\prime}\right|\right)\right\rceil$
            foreach $c\in\left\{1,\ldots,l\right\}$ do
                  $\left(s,v\right)\leftarrow\mathrm{choice}\left(D\right)$
                  remove $\left(s,v\right)$ from $D$
                  add $\left(s,v\right)$ in $D^{\prime}$
                  if $\left|D\right|=0$ then
                        $D\leftarrow D^{\prime}$
                        $D^{\prime}\leftarrow\emptyset$
                  $j\leftarrow\mathrm{choice}\left(G\right)$
                  $G\leftarrow G\backslash\{j\}$
                  if $\left|G\right|=0$ then
                        $G\leftarrow\left\{1,\ldots,g\right\}$
                  add $\left(s,v\right)$ in $S_{j}$
      foreach $i\in\left\{1,\ldots,g\right\}$ do
            minimize $\sum_{(s,v)\in S_{i}}\left(f_{\theta}(s)-v\right)^{2}$
Algorithm 8 Stochastic gradient descent with Stratified experience replay algorithm ($B$ is the batch size, $\mu$ is the memory size, $\delta$ is the duplication factor, $M$ is the memory buffer: a global variable initialized by an empty queue ; choice($E$): uniformly randomly chooses an element of $E$ ; see Table 1 for the definitions of the other symbols)
Remark 2.
We have not been able to evaluate the advantages or disadvantages of this variant with respect to the basic experience replay: the performances, we obtained during an experiment on this subject (not presented in this document), are so close that we would need increase the number of repetitions by 10 or maybe even 100 to be able to separate them, which is currently complicated or even impossible. We are not claiming that this algorithm is better than the classic replay experience, nor vice versa.
@@ 3.3 Common Experiment Protocol
The experiments of several sections share the same protocol. It is presented in this section. The protocol is used to compare different variants of reinforcement learning algorithms. A variant corresponds to a certain combination of elementary algorithms. More specifically, a combination consists of the association of a search algorithm (iterative deepening alpha-beta (with move ordering), MCTS (UCT with $c=0.4$ as exploration constant), $\mathrm{UBFM}$, …), of an action selection method ($\epsilon$-greedy distribution (used by default), softmax distribution, …), a terminal evaluation function $f_{\mathrm{t}}$ (the classic game gain (used by default), …), and a procedure for selecting the data to be learned (root learning, tree learning, or terminal learning). The protocol consists in carrying out a reinforcement learning process of $48$ hours for each variant (per repetition). At several stages of the learning process, each combination is evaluated. This evaluation consists in matches played by using the adaptive evaluation function generated by the evaluated combination. Each variant is thus characterized by a winning percentage at each stage of the reinforcement learning process. More formally, we denote by $f_{\theta_{h}}^{c}$ the evaluation generated by the combination $c$ at the hour $h$. Each combination is evaluated every hour by a winning percentage. The winning percentage of a combination $c$ at a hour $h\leq 48$ is computed from matches by using the minimax search at depth $1$ with $f_{\theta_{h}}^{c}$ as evaluation function confronting the following opponents: the minimax search at depth $1$ with $f_{\theta_{48}}^{c^{\prime}}$ as evaluation for all studied combination $c^{\prime}$ (there is one match in first player and another in second player per pair of combination).
@@
In addition to the curve, the different variants are also compared in relation to their final winning percentage, i.e. at the end of the learning process. Unlike the evolution of winning percentages, in the comparison of final performances, each evaluation $f_{\theta_{48}}^{c}$ confronts each other evaluation $f_{\theta_{48}}^{c^{\prime}}$ from all repetitions. In other words, this experiment consists in performing an all-play-all tournament with all the evaluation functions generated during the different repetitions. The presented winning percentage of a combination is still the average over the repetitions. The matches are also made by using minimax at depth $1$. These percentages are shown in tables.
@@
All the experiments based on this protocol involving MCTS were also performed with $c=\sqrt{2}$ as exploration constant. The results are similar.